The sum of four consecutive even numbers A, B, C, and D is 180. What is the sum of the next four consecutive even numbers?
(1) 214 (2) 212 (3) 196 (4) 204 (5) none of these

Answer by ptaylor(2198)

Let x=first integer
Then x+2=next integer
x+4=third integer
x+6=fourth consecutive integer
x+x+2+x+4+x+6=180
4x=180-12=168
x=42---first
x+42=42+2=44---2nd
44+2=46---3rd
46+2=48----4th
CK
42+44+46+48=180
180=180
Next 4 consecutive even integers then would be:
50+52+54+56=212
Hope this helps---ptaylor